# Strong convergence of fully discrete finite element schemes for the stochastic semilinear generalized Benjamin-Bona-Mahony equation driven by additive Wiener noise

### v1

@@ 6. Conclusion
In this work, we have analyzed the semi-discrete finite element approximation and the fully discrete scheme for the stochastic semilinear generalized BBM equation. Our results establish strong convergence rates with respect to both spatial and temporal discretization parameters.
For future research, we plan to extend the analysis to the stochastic semilinear generalized BBM-Burgers equation, obtained by incorporating a Burgers-type nonlinear term of the form $\nabla\cdot f(u)$. Another direction of interest is the study of weak convergence for the numerical approximation of the stochastic semilinear generalized BBM equation. Furthermore, we intend to investigate the strong convergence behavior of the scheme when the driving noise is a multiplicative $Q$-Wiener process and Lévy process.

### v2

@@ 6. Conclusion
In this work, we have analyzed the semi-discrete finite element approximation and the fully discrete scheme for the fourth-order stochastic pseudo-parabolic equation. Our results establish strong convergence rates for both spatial and temporal discretization parameters.
An interesting direction for future research is the investigation of weak convergence for the numerical approximation of fourth-order stochastic pseudo-parabolic equations. Furthermore, we intend to investigate the strong convergence behavior of the scheme when the driving noise is a multiplicative $Q$-Wiener process and Lévy process.
@@
Declarations:
Ethics approval and consent to participate:
Not Applicable.
Consent for Publication: Not Applicable.
Funding: The authors declare that there was no funding for this work.
Data availability: No data were used during the study.
Author contributions: Suprio Bhar has supervised the project and reviewed the manuscript. Mrinmay Biswas has selected the problem, reviewed the manuscript and contributed in the numerical simulation parts. Managala Prasad has done most of the technical part such as writing proofs and calculations, and has written the main manuscript.